NicktheBlade16:

Solve log base 2 of one eighth. −3 3 negative one third one third

NicktheBlade16:

Alexis1415:

Ok now this one is -3 because, \[\log _{2} (1/8) = \log _{2}1-\log _{2}8\] \[0-\log _{2}8\] Then we change the base to 10 So, \[(\log 8) /(\log 2)\] Which gives us -3

NicktheBlade16:

Alright thank you.

Alexis1415:

@nicktheblade16 wrote:
Alright thank you.
ofc